# supplementary_information

### v1

@@ Abstract
Performance comparisons are fundamental in medical imaging Artificial Intelligence (AI) research, often driving claims of superiority based on relative improvements in common performance metrics. However, such claims frequently rely solely on empirical mean performance. In this paper, we investigate whether newly proposed methods genuinely outperform the state of the art by analyzing a representative cohort of medical imaging papers. We quantify the probability of false claims based on a Bayesian approach that leverages reported results alongside empirically estimated model congruence to estimate whether the relative ranking of methods is likely to have occurred by chance. According to our results, the majority ($>$80%) of papers claims outperformance when introducing a new method. Our analysis further revealed a high probability ($>$5%) of false outperformance claims in 86% of classification papers and 53% of segmentation papers. These findings highlight a critical flaw in current benchmarking practices: claims of outperformance in medical imaging AI are frequently unsubstantiated, posing a risk of misdirecting future research efforts.
@@ Results
Figure 3: Claims of outperformance in medical classification (a) and segmentation (b) papers are not well-justified. The majority of papers ($>$ 80%) claims outperformance and highlights the best performance metric values in bold. A low portion ($<$15%) backs up the claims by further analyses, such as statistical testing.
@@
Our analysis revealed a high probability ($>$5%) of false claims in the majority of medical imaging AI publications. Specifically, 86% of classification papers (Figure 4a) and 53% of segmentation papers (Figure 4b) exhibited a false claim probability exceeding this threshold, when computing the results based on the median congruence values. Further, for 58% of classification papers (Figure 5a) and 25% of segmentation papers (Figure 5b) we even found an extremely high probability of false claims of over 30%.
Figure 4: Our analysis reveals a high likelihood for false outperformance claims in medical image classification and segmentation. (a) For more than 86% of all MICCAI 2023 classification papers, we estimated a high probability ($>$5%) of false claims of outperformance of their first-ranked method over the second-ranked one. The dots correspond to the reported difference in Accuracy between the two top-ranked methods of the MICCAI 2023 classification papers and their respective test set sizes. (b) For more than 53% of all MICCAI 2023 segmentation papers, we estimated a high probability ($>$5%) of false claims of outperformance of their first-ranked method over the second-ranked one. The dots correspond to the reported difference in DSC between the two top-ranked methods of the MICCAI 2023 segmentation papers and their respective test set sizes. For both figures, the x-axis, representing the test set size, is displayed on a logarithmic scale.
@@
Figure 6: Substantiating outperformance requires substantially larger sample sizes. Probability of false claims as a function of test set size and metric difference for (a) classification papers and (b) segmentation papers. Each cell displays the median probability based on assumed median model congruence, with values for the 25th and 75th percentile model congruence in parentheses. Color coding: red ($>$10%), orange (5–10%), green ($<$5%). The left gray-highlighted cell represents observed MICCAI 2023 values, while the right gray-highlighted cell indicates test set sizes leading to $<$5% false claim probability for a fixed range of performance deltas.
@@ Discussion
False outperformance claims in medical imaging AI research often stem from factors that give models an unfair advantage during evaluation, making them seem more effective than they actually are. Two major contributing factors for this are model overfitting and—more generally—data leakage, both of which lead to overly optimistic performance estimates and poor generalization to new data. Model overfitting often occurs when authors repeatedly test different models on identical test sets and then report on the performance of the best model. Data leakage is a more general concept that refers to a situation where information that should not be available during the model training phase is inadvertently included, leading to overly optimistic performance estimates or inaccurate predictions. This occurs because the model learns patterns or relationships from the “leaked” data. Common types of leakage include target leakage, train-test-contamination (e.g., when preprocessing steps like normalization or feature engineering are performed on the entire dataset before splitting it into training and test sets) and temporal leakage in time-series data [1]. Small sample sizes make these problems worse as they undermine statistical power of model evaluation in AI [2]. Publication bias, with corresponding incentives and selective reporting, are known to make statistical problems worse [3]. Our results reveal the pervasiveness of these failures in medical image AI [2].
@@
A limitation of our study could be seen in the fact that we did not have access to the true congruence parameters. To estimate the probability of false claims, we relied on external datasets to approximate these key parameters for both segmentation and classification. For segmentation, we used the Medical Segmentation Decathlon (MSD) dataset to calculate the correlation coefficient of the test set DSC and to approximate unreported standard deviation values in the analyzed MICCAI 2023 papers [4]. For classification, we extrapolated information on the assumed congruence from a comprehensive private dataset. While these choices were driven by pragmatic considerations, and generalizability could be a potential limitation, our sensitivity analyses confirm that the methodology remains robust across varying parameter assumptions.
@@ Methods
[…]
@@
  • $t_{n-1}$ is the quantile of the Student distribution with $n-1$ degrees of freedom.
@@
To get an empirical estimate for model congruence for classification tasks, we selected a diverse set of 12 medical imaging tasks from [5, 6] covering eight different modalities, namely fundus photography , confocal laser endomicroscopy [7] , computed tomography [8], dermatoscopy [9, 10, 11], gastroscopy and colonoscopy [12], sonography [13], and X-ray [14, 15]. Our task selection aimed to cover sufficient variety in modality, number of samples and class imbalance. The median test set size was 304 images. We then applied a total of 5 [16, 17, 18, 19, 20] different architectures to all tasks and calculated model congruence values for Accuracy from which we retained the median (0.67), 25th percentile (0.47), and 75th percentile (0.83). Since model congruence values for Accuracy are constrained by lower and upper bounds (see Supplementary information, Subsection C) we adapted the extrapolation of these derived congruence values to the eligible MICCAI 2023 classification papers. The procedure involved first assigning an assumed model congruence value (either the median, 25th percentile, or 75th percentile derived from our private dataset) to each paper. We then assessed whether these extrapolated values fell within the acceptable bounds defined by the characteristics of each MICCAI 2023 classification paper. If an extrapolated value exceeded those bounds, it was adjusted by replacing it with either the lower or upper bound, selecting whichever value had the smallest difference from the original extrapolated value.
@@
Additionally, to apply the formula for estimating the probability of false claims in segmentation papers, standard deviation (SD) values of the DSC for each of the two methods being compared were required. While the test set sample sizes (N) were generally accessible in the reported data, SD values were frequently missing (see Figure S1). To address this, we imputed the missing standard deviation values following the approach described in [21]. Similarly to what was done for congruence, we assessed the sensitivity of the probability of false claims to this imputation. To that purpose, we computed the prediction interval (PI) for the SD imputation and repeated our analysis with the first quartile (Q1) and third quartile (Q3) of the PI [21, 22]. Results show that our conclusions are robust with respect to the imputation of SD (Figure S4).

### v2

# False Promises in Medical Imaging AI? Assessing Validity of Outperformance Claims

@@ Abstract
Performance comparisons are fundamental in medical imaging Artificial Intelligence (AI) research, often driving claims of superiority based on relative improvements in common performance metrics. However, such claims frequently rely solely on empirical mean performance. In this paper, we investigate whether newly proposed methods genuinely outperform the state of the art by analyzing a representative cohort of medical imaging papers. We quantify the probability of false claims based on a Bayesian approach that leverages reported results alongside empirically estimated model congruence to estimate whether the relative ranking of methods is likely to have occurred by chance. According to our results, the majority (>>80%) of papers claims outperformance when introducing a new method. Our analysis further revealed a high probability (>>5%) of false outperformance claims in 86% of classification papers and 53% of segmentation papers. These findings highlight a critical flaw in current benchmarking practices: claims of outperformance in medical imaging AI are frequently unsubstantiated, posing a risk of misdirecting future research efforts.
@@ Results
Figure 3: Claims of outperformance in medical classification (a) and segmentation (b) papers are not well-justified. The majority of papers (>> 80%) claims outperformance and highlights the best performance metric values in bold. A low portion (<<15%) backs up the claims by further analyses, such as statistical testing.
@@
Our analysis revealed a high probability (>>5%) of false claims in the majority of medical imaging AI publications. Specifically, 86% of classification papers (Figure 4a) and 53% of segmentation papers (Figure 4b) exhibited a false claim probability exceeding this threshold, when computing the results based on the median congruence values. Further, for 58% of classification papers (Figure 5a) and 25% of segmentation papers (Figure 5b) we even found an extremely high probability of false claims of over 30%.
Figure 4: Our analysis reveals a high likelihood for false outperformance claims in medical image classification and segmentation. (a) For more than 86% of all MICCAI 2023 classification papers, we estimated a high probability (>>5%) of false claims of outperformance of their first-ranked method over the second-ranked one. The dots correspond to the reported difference in Accuracy between the two top-ranked methods of the MICCAI 2023 classification papers and their respective test set sizes. (b) For more than 53% of all MICCAI 2023 segmentation papers, we estimated a high probability (>>5%) of false claims of outperformance of their first-ranked method over the second-ranked one. The dots correspond to the reported difference in DSC between the two top-ranked methods of the MICCAI 2023 segmentation papers and their respective test set sizes. For both figures, the x-axis, representing the test set size, is displayed on a logarithmic scale.
@@
Figure 6: Substantiating outperformance requires substantially larger sample sizes. Probability of false claims as a function of test set size and metric difference for (a) classification papers and (b) segmentation papers. Each cell displays the median probability based on assumed median model congruence, with values for the 25th and 75th percentile model congruence in parentheses. Color coding: red (>>10%), orange (5–10%), green (<<5%). The left gray-highlighted cell represents observed MICCAI 2023 values, while the right gray-highlighted cell indicates test set sizes leading to <<5% false claim probability for a fixed range of performance deltas.
@@ Discussion
False outperformance claims in medical imaging AI research often stem from factors that give models an unfair advantage during evaluation, making them seem more effective than they actually are. Two major contributing factors for this are model overfitting and—more generally—data leakage, both of which lead to overly optimistic performance estimates and poor generalization to new data. Model overfitting often occurs when authors repeatedly test different models on identical test sets and then report on the performance of the best model. Data leakage is a more general concept that refers to a situation where information that should not be available during the model training phase is inadvertently included, leading to overly optimistic performance estimates or inaccurate predictions. This occurs because the model learns patterns or relationships from the “leaked” data. Common types of leakage include target leakage, train-test-contamination (e.g., when preprocessing steps like normalization or feature engineering are performed on the entire dataset before splitting it into training and test sets) and temporal leakage in time-series data [kapoor2023leakage]. Small sample sizes make these problems worse as they undermine statistical power of model evaluation in AI [varoquaux2018cross]. Publication bias, with corresponding incentives and selective reporting, are known to make statistical problems worse [ioannidis2005most]. Our results reveal the pervasiveness of these failures in medical image AI [varoquaux2018cross].
@@
A limitation of our study could be seen in the fact that we did not have access to the true congruence parameters. To estimate the probability of false claims, we relied on external datasets to approximate these key parameters for both segmentation and classification. For segmentation, we used the Medical Segmentation Decathlon (MSD) dataset to calculate the correlation coefficient of the test set DSC and to approximate unreported standard deviation values in the analyzed MICCAI 2023 papers [antonelli2022medical]. For classification, we extrapolated information on the assumed congruence from a comprehensive private dataset. While these choices were driven by pragmatic considerations, and generalizability could be a potential limitation, our sensitivity analyses confirm that the methodology remains robust across varying parameter assumptions.
@@ Methods
$P(\theta_{A}\leq\theta_{B}\mid\textbf{reported results})=P(p_{A}\leq p_{B}\mid\widehat{p}_{A},\widehat{p}_{B})=P(p_{1}\leq p_{2}\mid\widehat{p}_{A},\widehat{p}_{B})=\int_{0}^{1}\int_{0}^{p_{2}}p(p_{1},p_{2}|\widehat{p}_{A},\widehat{p}_{B})dp_{1}dp_{2}$ (1)
@@
  • $t_{n-1}$ is the quantile of the Student distribution with n1n-1 degrees of freedom.
@@
To get an empirical estimate for model congruence for classification tasks, we selected a diverse set of 12 medical imaging tasks from [godau2024fingerprinting, godau2024mml] covering eight different modalities, namely fundus photography , confocal laser endomicroscopy [aidae2024challenge] , computed tomography [yang2020covidct], dermatoscopy [kawahara2018seven, isic2020challenge, rotemberg2021melanoma], gastroscopy and colonoscopy [borgli2020hyperkvasir], sonography [aldhabyani2020breast], and X-ray [rajpurkar2017mura, irvin2019chexpert]. Our task selection aimed to cover sufficient variety in modality, number of samples and class imbalance. The median test set size was 304 images. We then applied a total of 5 [yu2023metaformer, wu2022tinyvit, liu2021swin, dai2021coatnet, yalniz2019billion] different architectures to all tasks and calculated model congruence values for Accuracy from which we retained the median (0.67), 25th percentile (0.47), and 75th percentile (0.83). Since model congruence values for Accuracy are constrained by lower and upper bounds (see Supplementary Information, Subsection C) we adapted the extrapolation of these derived congruence values to the eligible MICCAI 2023 classification papers. The procedure involved first assigning an assumed model congruence value (either the median, 25th percentile, or 75th percentile derived from our private dataset) to each paper. We then assessed whether these extrapolated values fell within the acceptable bounds defined by the characteristics of each MICCAI 2023 classification paper. If an extrapolated value exceeded those bounds, it was adjusted by replacing it with either the lower or upper bound, selecting whichever value had the smallest difference from the original extrapolated value.
@@
Additionally, to apply the formula for estimating the probability of false claims in segmentation papers, standard deviation (SD) values of the DSC for each of the two methods being compared were required. While the test set sample sizes (N) were generally accessible in the reported data, SD values were frequently missing (see Figure S1). To address this, we imputed the missing standard deviation values following the approach described in [10.1007/978-3-031-72117-5_12]. Similarly to what was done for congruence, we assessed the sensitivity of the probability of false claims to this imputation. To that purpose, we computed the prediction interval (PI) for the SD imputation and repeated our analysis with the first quartile (Q1) and third quartile (Q3) of the PI [10.1007/978-3-031-72117-5_12, wasef2016derivation]. Results show that our conclusions are robust with respect to the imputation of SD (Figure S4).